# ScaleRAFT: Cross-Scale Recurrent All-Pairs Field Transforms for 3D Motion Estimation

### v1

@@ Abstract
In this paper, we study the problem of estimating the 3D motion of dense pixels from continuous image pairs. Most previous methods are based on mature optical flow baselines and depth values, projecting the 2D motion on pixel planes into 3D space, and further optimizing the results by combining depth-motion-branch and other sub-modules. This stacked framework cannot leverage the complementarity between optical flow and other modules nor escape the dependence on accurate depth information. To address the above challenges, we propose a normalized scene flow framework, ScaleRAFT, based on cross-scale matching. Its core feature is directly matching objects between two frames in 3D scale space, i.e. matching features at the correct location and scale. Unlike previous methods, ScaleRAFT integrates optical flow and deep motion estimation into a unified architecture, allowing the optical flow pipeline and deep motion estimation to promote each other mutually. Moreover, ScaleRAFT estimates motion in the depth direction based on feature matching, breaking away from the dependence on accurate depth information. Experimentally, our method has achieved the best foreground performance so far in motion estimation tasks in driving scenarios, and has significantly improved various downstream 3D tasks.
@@ I Introduction
Three-dimensional flow aims to estimate the motion trend of pixel points in a pair of images in 3D space, which is numerically composed of optical flow and motion in the depth directions. It is a long-standing and important problem in computer vision[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
Figure 1: Tradition matching vs cross-scale matching (ScaleRAFT). From top to bottom are: (a,b) two consecutive frames of input images, (c,d) optical flow error maps based on tradition matching and cross-scale matching, respectively, (e) motion in the depth direction based on ScaleRAFT, (f) The collision time of the foreground targets predicted based on ScaleRAFT. Notice that in (c,d), the foreground error of the baseline method is greatly reduced (96.8$\rightarrow$61.29), thanks to cross-scale matching reducing the impact of scale changes on optical flow matching. In addition, CSCV also provides key clues for deep motion, and (e, f) demonstrates its application in 3D visual tasks.
Previous Works. Most advanced 3D flow methods are based on mature optical flow pipelines [13, 14, 15, 16, 17]. They project the pixel flow results estimated by the optical flow pipeline into 3D space through depth values, and add sub-modules such as deep motion branches or rigid optimization to enhance the results further. Although they have achieved remarkable results, their framework highly relies on the optical flow pipeline and input depth values, so defects in the optical flow module or depth value will directly damage the overall performance. There is also a type of 3D flow method[18, 19] based on monocular images; They regress motion-in-depth (MID) from optical flow or image pairs, resulting in sub-precise and low robustness results. Moreover, they separated the estimated pipelines of optical flow and MID, without utilizing the complementarity between pipelines.
In addition to the problems of the 3D flow framework itself, the optical flow pipeline as its cornerstone also has inherent flaws that have been ignored for a long time. Most optical flow methods[20, 21, 22, 23, 24] rely on feature-matching techniques, correlating objects in two frames by comparing the correlation between features. However, we found that feature matching in optical flow is limited to the original scale. When an object moves along the camera’s Z-axis (depth direction), its scale change often leads to feature-matching failure. Fig. 3 shows a typical example where, except for the occluded area, the principal errors of the optical flow algorithm are concentrated in the area where scale changes occur, indicating that scale changes are one of the essential reasons for optical flow errors.
Inspiration. The inspiration for solving the above problem comes from the SIFT[25] matching algorithm proposed by Lowe et al. in 2004. They also recognized the impact of scale factors on feature point matching and constructed a scale pyramid to extract scale-invariant features, ultimately achieving stable cross-scale feature point matching. Further, in 2005, Mikolajczyk[26] successfully estimated the scale changes of the clock tower captured at different focal lengths using similar scale selection techniques. Through the successful experience of predecessors, we recognize that features have good stability and separability in the scale dimension. Therefore, we attempt to extend optical flow matching from a 2D pixel plane to a 3D scale space, match features at the correct position and scale, solve the impact of scale changes on optical flow matching, and provide more depth clues by estimating the scale change.
@@
Figure 3: Box plot of optical flow error distribution in different scale change regions. We measure the scale change of different regions based on scale change rate $s_{p}$, which means the size change percentage of objects between two frames. We pre-trained the optical flow network RAFT on Flyingthings3D and evaluated it on the test set. During the evaluation, $s_{p}$ is divided into three intervals: $(0,2.5\%),(2.5\%,32.5\%)$, and $(32.5\%,+\infty)$, corresponding to small(S), medium(M), and large(L) scale change regions. The subgraph title provides the percentage of each region in the test dataset. Only non-occluded pixels have been included in the evaluation. Results indicate that the error of the optical flow baseline is concentrated in regions with significant scale changes, which is consistent with our motivational analysis and Lowe et al.’s judgment[25].
Our Method. We propose an enhanced cross-scale normalized scene flow (NSF) search framework based on RAFT (ScaleRAFT). The core feature is to replace the original single scale correlation volume (CV) with cross-scale correlation volume (CSCV), to achieve cross-scale feature matching by calculating the correlation between different position and scale feature pairs. As shown in Fig. 2, CSCV matches objects with different motion velocities (depth direction) at different scales, achieving more stable feature matching. The matching behavior in the scale dimension also directly provides the scaling factor of each object between two frames. Based on previous work[18], we can naturally obtain the corresponding depth change rate, which helps the network alleviate dependence on regression and depth values. We demonstrate in Fig. 1 that ScaleRAFT can robustly estimate 3D motion and extend it to downstream 3D tasks.
@@
Expect scale changes, there are typical occlusion and overfitting issues in three-dimensional flow. We have improved the commonly used truth supervision mode and proposed a training pipeline that combines self-supervision and truth. Specifically, we iteratively generate flight foreground with random textures and shapes to cover the original image pairs, and calculate the 3D flow truth of the flight foreground to replace the original position labels. Previous works often use solid coloured blocks to cover the source image to create occlusion actively. However, this behaviour encouraged the network to rely overly on regression to estimate optical flow, resulting in poor generalization. Our hybrid training pipeline actively creates occlusion while forcing the network to learn the 3D flow of randomly moving foreground through texture matching, ultimately balancing occlusion prediction and texture matching learning.
To broaden the application scope of our framework, we added a rigid background optimization module (RBO) after the basic ScaleRAFT in this paper. Based on depth values and semantic classification, RBO further optimized the background part of the 3D flow, upgrading the original normalized scene flow method to the RGBD-based scene flow method (called SacleRAFT-RBO). Moreover, to further verify the rationality and universality of the cross-scale matching approach, we ported the core module CSCV to the advanced scene flow method RAFT3D (called SacleRAFT3D) and directly replaced the original CV correlation volume with CSCV, changed the calculation method of deep motion from regression to scale matching. Compared with RAFT3D, ScaleRAFT3D changes the direct regression prediction mode, providing more reliable clues for depth motion prediction.
Significant performance improvement validates our motivation. On KITTI[27], ScaleRAFT far surpasses all NSF methods, reducing SF-all to 6.12%. As a monocular method, ScaleRAFT even outperforms the method using additional depth information (LiDAR, RGB-D) in foreground metrics, refreshing the optimal performance of foreground scene flow and foreground optical flow. ScaleRAFT3D also achieved significant performance improvement compared to baseline RAFT3D, with SF-all reduced from 5.77% to 4.74%, proving that our cross-scale matching technique has a certain degree of universality. In addition, although not specifically designed for RGBD, our ScaleRAFT-RBO still achieves leading performance with SF-all only at 4.20%. On Flyingthings3D[28], we followed the training and testing process of RAFT3D. The end-to-end 3D error of ScaleRAFT3D was reduced from 6.4cm to 4.9cm compared to the baseline model. Our method also has strong generalization and robustness. We demonstrated the excellent performance of ScaleRAFT in the field driving scenes without fine-tuning and daily scenes with non-rigid objects in the Supplementary Material Sec. III.
  • We propose a simple yet powerful normalized scene flow framework, ScaleRAFT, which unifies optical flow and depth motion estimation into one model, achieving more robust 3D flow estimation.
  • We propose a self-supervised truth mixing training pipeline that balances the learning of occlusion regression and texture matching.
  • We built two different directions of extension ScaleRAFT-RBO and ScaleRAFT3D based on the basic framework ScaleRAFT, expanding the application scope of the ScaleRAFT framework and proving the universality of cross-scale matching ideas.
  • We test our methods on various 3D flow tasks and generally achieve significant improvements over previous state-of-the-art methods.
Differences from The Preliminary Version. A preliminary version of this work, Scale-flow[29], has been accepted by ACM MM 2022. Compared to the preliminary version, our expanded paper has added several new contents: (i) We introduce a self-supervised and truth mixed training pipeline that helps the network better balance the learning between occlusion fitting and texture matching, alleviate potential overfitting problems during training, and better predict motion prospects. (ii) We have extended the original monocular method to an RGBD version by adding a background rigid optimization module, expanding the application scope of our method. (iii) We have extended the core module CSCV to other advanced RGBD scene flow methods, further verifying the effectiveness of cross-scale matching ideas. (iv) We also demonstrated more ablation studies, particularly analyzing the effectiveness of mixed training pipelines. (v) We validated the effectiveness of our proposed method and two extensions on the official KITTI testing benchmark. [^1] https://www.cvlibs.net/datasets/kitti/eval_scene_flow.php (vi) We further validated our method’s scene flow performance on the Flyingthings3D dataset.
We have released a demo program [^2] https://github.com/HanLingsgjk/CSCV for 3D motion estimation in daily scenes and evaluation corresponding to the key results in the article. Welcome to check.
@@ II-A Optical Flow
Early methods[30, 31, 32, 33] viewed optical flow as an optimization problem for minimizing energy. Based on the regularization term and the constant brightness assumption, the optical flow results are optimized iteratively. Some recent excellent works have inherited their ideas[20, 34, 13], gradually updating the optical flow field through correlation matching and warping techniques. Compared with the previous methods [35, 21, 36, 37, 38, 39, 40, 41, 42, 43] of directly using convolutional networks to regress optical flow, the method based on correlation matching has a clearer theoretical background and more robust results. In this work, inspired by the successful experience of correlation matching, we extended this matching pattern to the scale dimension based on CSCV, improving the scale robustness of the optical flow baseline model.
It is worth mentioning that although some previous optical flow works[44, 45, 20, 46, 47, 48, 49] also used the idea of scale, our method differs fundamentally from theirs. Previous works are often based on multi-scale technology to increase the network receptive field, improve the inference speed, and strengthen the loss function. However, it is still image-to-image matching in essence. CSCV introduces a pattern for optical flow matching in scale space, a brand-new attempt.
@@ II-B Scene Flow
Methods Based on Point Cloud. Point cloud based methods[52, 53, 54, 55, 56, 57, 58, 59, 60, 61] capture 3D point clouds of the surrounding environment through an active laser or stereo camera, and establish the corresponding relationship between two frame point clouds to estimate 3D motion. However, this kind of method is limited by the sparse results and limited range, which is challenging to apply in more common video scenes. There is also a point cloud method based on the fusion framework, Camliflow[14]. By highly fusing the point cloud and the optical flow pipeline, the two pipelines can obtain complementary results at the same time. Camliflow achieves end-to-end scene flow estimation, but it still has the problem of depth information dependence.
@@
II-C Motion-in-depth (MID)
MID describes the motion in the depth direction through the scale change of depth. It has many practical application scenarios, such as time-to-collision (TTC) estimation, LiDAR scene flow, scene flow, Etc. MID is also the core indicator of our ScaleRAFT.
Early TTC works[62, 63, 64] estimated the MID by tracking the motion trajectory of the interest points and building a motion model. Because the distribution of the interest points is random, this kind of method often makes it hard to pay attention to the primary target, yielding only sparse and low-precision results. Binary TTC[19] is a successful TTC scheme proposed recently, which can regress a dense MID field segmentally. Binary TTC sends the scaled image to the encoder to directly return the MID result of a certain range, and the specific range depends on the scaling factor. Therefore, a complete MID estimation requires multiple regressions to cover the complete MID interval. Unlike them, ScaleRAFT only needs one calculation to get complete and dense MID results.
Yang & Ramanan[18] proposed the only monocular normalized scene flow method. They obtained dense MID results based on optical flow step-wise regression optical expansion and MID. However, the same as previous TTC work, they separated optical flow and MID estimation, resulting in the depth motion estimation result dependent on the optical flow result. Unlike them, ScaleRAFT integrates optical flow and depth motion estimation into a harmonious task based on CSCV, making information blend and help each other. In addition, the cross-scale matching mechanism can also help the optical flow module more accurately estimate the objects with scale changes.
II-D Cross-scale Matching
Cross-scale matching aims to establish the correspondence between objects in different frames[65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75], and the scales of objects in these frames often have large changes, which is a difficult problem faced by current optical flow methods. In the early works, SIFT[25] obtains scale-invariant features in the scale space through hand-designed feature extraction operators, and then matches objects of different scales based on these stable features. In addition to being used for matching, cross-scale matching technology can also be used to estimate the scale changes of objects between frames. As early as 2005, Mikolajczyk[26] used Lindeberg’s scale selection method[76] to successfully estimate the optical expansion of clock towers captured at different focal lengths. Inspired by the above works, CSCV performs optical flow matching in scale space and estimates a dense scale change field based on the cross-scale matching results.
Figure 5: Overview of ScaleRAFT. First, the frame2 is multi-scaled by bilinear interpolation, and fed to the feature extractor with the frame1 to obtain multi-scale features. Then a cross-scale correlation volume (CSCV) is established based on the multi-scale feature pairs, from which the correlation features about the current field can be sampled. Finally, the extracted correlation features are sent to the GRU optimizer to iteratively optimize an optical flow field and a scale change field. The MID result can be converted according to the scale change field by Eq.19.
III OUR APPROACH
Recurrent All-pairs Field Transforms (RAFT)[20] is the state-of-the-art optical flow baseline[77], which constructs a 4D correlation volume over the all-pairs pixel and iteratively optimizes the optical flow field through the GRU optimizer. In this section, we use a cross-scale matching module (CSCV) to reconstruct the RAFT baseline and propose the new monocular 3D flow method called ScaleRAFT. The 3D flow estimated by ScaleRAFT consists of a 2D optical flow field and a 1D motion-in-depth (MID) field, where the MID is obtained by converting the scale change field.
The framework structure of ScaleRAFT is shown in Fig. 5. CSCV replaces the original CV module in RAFT, and an additional optical expansion output head is added after the GPU module to predict the scale change field $f_{3}$ required for CSCV sampling. Combining with previous work[18], we convert the scale change field $f_{3}$ into motion-in-depth (MID), so that ScaleRAFT can be widely used in various monocular 3D flow tasks.
It is worth mentioning that ScaleRAFT is an end-to-end monocular 3D flow method. It integrates optical flow and depth motion prediction into a unified and harmonious matching framework through CSCV. We demonstrate in experiments that ScaleRAFT has impressive performance improvements and better robustness compared to previous monocular 3D flow methods. In the following parts, we will introduce in sequence: The construction and training process of ScaleRAFT (III-A,III-B,III-C,III-D); How to convert the results of network estimation into a 3D motion field (III-E); RGB-D extension based on rigid background optimization (III-F); Extend CSCV to other RGBD methods (III-G).

### v2

# ScaleFlow++: Robust and Accurate Estimation of 3D Motion from Video

@@ Abstract
Perceiving and understanding 3D motion is a core technology in fields such as autonomous driving, robots, and motion prediction. This paper proposes a 3D motion perception method called ScaleFlow++ that is easy to generalize. With just a pair of RGB images, ScaleFlow++ can robustly estimate optical flow and motion-in-depth (MID). Most existing methods directly regress MID from two RGB frames or optical flow, resulting in inaccurate and unstable results. Our key insight is cross-scale matching, which extracts deep motion clues by matching objects in pairs of images at different scales. Unlike previous methods, ScaleFlow++ integrates optical flow and MID estimation into a unified architecture, estimating optical flow and MID end-to-end based on feature matching. Moreover, we also proposed modules such as global initialization network, global iterative optimizer, and hybrid training pipeline to integrate global motion information, reduce the number of iterations, and prevent overfitting during training. On KITTI, ScaleFlow++ achieved the best monocular scene flow estimation performance, reducing SF-all from 6.21 to 5.79. The evaluation of MID even surpasses RGBD-based methods. In addition, ScaleFlow++ has achieved stunning zero-shot generalization performance in both rigid and nonrigid scenes. Code is available at https://github.com/HanLingsgjk/CSCV.
@@ I Introduction
Three-dimensional motion estimation aims to estimate the motion trend of pixels in a pair of images in 3D space. It has essential applications in video understanding[1, 2], autonomous driving[3, 4, 5, 6, 7, 8, 9], and autonomous robot navigation[10]. For example, in autonomous driving scenarios[11, 12], the car platform needs to obtain the 3D motion state of typical targets to make path decisions. Most existing platforms use Lidar to obtain 3D point clouds of scenes and construct corresponding relationships between two frames of 3D point clouds to estimate 3D motion. However, due to limited range and sparse results, this active depth perception scheme cannot effectively handle objects at a distance and abnormally reflective surfaces. In addition, the expensive price and maintenance costs further limit the broader application of Lidar. In this paper, we consider using a monocular camera to estimate 3D motion in dynamic scenes (commonly called normalized scene flow), which is a more stable and easy-to-maintain sensor that is not limited by limited range and sparse results.
Challenge: Normalized scene flow (NSF) consists of optical flow and motion-in-depth. Monocular cameras can robustly estimate 2D optical flow and normalized camera motion trajectories by constructing pixel-by-pixel feature associations between two frames[13, 14]. However, monocular cameras cannot provide real scene depth[15]. In dynamic scenes, 2D optical flow is a function of scene speed, camera motion trajectory, and scene depth. Given only the optical flow and camera trajectory, solving these three parts is an under-constrained problem.
Figure 1: Zero-shot Generalization in Real Scenes. Only needing to input two consecutive frames of images, ScaleFlow++ can robustly estimate dense 3D motion fields. The lower left corner of the image is the color-coded optical flow map, and the lower right corner is the motion-in-depth (MID). MID provides additional depth cues in the Z-axis direction, such as the dancer’s left leg moving away from the camera and the right leg moving closer to the camera.
Previous work: In early time-to-collision (TTC) work[16, 17, 18], people found that there is a close correlation between the scale change of the object and the depth motion. They estimated the scale change of a specific image block through algorithms such as SIFT[19], and converted the scale change into sparse depth motion through the transformer formula. Recent works[20, 21] regress dense scale changes from optical flow fields and RGB images through deep networks. However, their over-reliance on regression leads to lower accuracy and unstable results. Moreover, they separate the calculation of optical flow and motion-in-depth, requiring the network to be trained in stages rather than end-to-end training.
In addition to the regression dependency dilemma faced by motion-in-depth estimation, existing optical flow pipelines are also plagued by scale issue. Most optical flow methods[22, 23, 24, 25, 26] rely on feature-matching techniques to associate objects in two frames by comparing the correlations between features. However, previous studies have pointed out that the commonly used CNN feature extractors only satisfy translation invariance and do not satisfy scale invariance[27]. When an object moves along the Z-axis (depth direction) of the camera, its scale change often leads to feature-matching failure. We demonstrate the significant impact of scale variation on feature matching in Fig. 2(c).
@@
Our method: Fortunately, previous work has already explored many issues related to scale. As early as 1998, Lin et al.[28]. proposed a scheme based on the Laplace operator to calculate the maximum response position in the scale pyramid to solve the problem of scale change in feature recognition. Even if the scale of the input image changes, the output response features are stable. Furthermore, in 2005, Mikolajczyk[29] successfully estimated the scale changes of the bell tower captured at different focal lengths using a similar approach. Therefore, inspired by the success of previous work, we propose a cross-scale correlation volume (CSCV) module. As shown in Fig. 2, CSCV extends the original optical flow matching from the 2D-pixel plane to the scale pyramid, matching the correct features at the correct position and scale, improving the previous dilemma of estimating MID overly dependent on regression.
We propose an enhanced monocular 3D flow estimation network ScaleFlow++ based on the successful optical flow framework RAFT[22]. Its core feature is to replace the original 4D correlation volume with CSCV, and estimate both optical flow and MID based on feature matching in a unified architecture. To enhance the information aggregation capability of the network, we propose a lightweight global iterative optimizer (GIR). In previous studies, most RAFT-style methods used the GRU optimizer. However, this simple single-scale optimizer can only provide a local view, which makes the optimizer overly focused on the correlation features obtained from local sampling and unable to coordinate global motion information over a larger scale range. On the contrary, GIR referred to Unet’s[30] successful experience and constructed an hourglass-shaped multi-scale residual convolutional network. In the encoding stage, it used the ConvNextV2[31] module with a super large convolution kernel and employed pyramid pooling[32] at the minimum scale to coordinate global information. The results indicate that using the GIR optimizer can significantly improve background performance and slightly enhance foreground performance while keeping the total time almost constant.
In addition to GIR, we also introduced a global initialization module, which is mainly due to the task alienation observed in RAFT-style methods at different iteration periods. In the early stages of iteration, the optimizer starts optimizing from zero, which requires the optimizer to make aggressive estimates. At the end of the iteration, the optimizer focuses more on fine-tuning. The difficulty and focus of tasks vary significantly at different stages of iteration. Therefore, in order to alleviate the training burden of task alienation on the optimizer, we used an additional optimizer to quickly initialize the entire motion field at a scale of 1/16, to alleviate the alienation problem in subsequent iterations and reduce the total number of iterations.
@@
During the training phase, we improved the commonly used truth supervision mode and proposed an enhanced training pipeline that combines self-supervised mixed truth. Specifically, we generate a flying foreground with random textures and shapes, iteratively overlay it on the original image pairs, and calculate the 3D stream truth of the flying foreground as a self-supervised training label. Previous works[22] often used solid color blocks to cover the source image in order to actively create occlusion. However, this behavior encourages the network to overly rely on regressing to estimate optical flow, resulting in poor generalization ability. Our hybrid training pipeline actively creates occlusions while forcing the network to learn 3D flow of randomly moving foreground through texture matching, ultimately balancing occlusion prediction and texture matching learning.
The significant performance improvement validates our motivation. In the KITTI[33] scene flow test set, ScaleFlow++ surpassed all NSF methods, reducing the core metric SF-all from 6.21% to 5.79%. In MID estimation, the minimum MID error was reduced from 42.84 to 38.44. In the Sintel[34] optical flow test set, ScaleFlow++ outperforms most similar methods with highly competitive performance. Moreover, as shown in Fig. 1, ScaleFlow++ also has strong generalization ability and robustness. In Sec. IV-G, we demonstrated the outstanding performance of ScaleFlow++ in unseen scenarios.
Differences from the preliminary version: This work is a substantial extension of our previous paper, Scale-flow[35], presented at the ACM MM 2022 conference. The preliminary work proposes a monocular 3D flow estimation method based on cross-scale matching, which alleviates the previous dilemma of relying solely on regression to calculate MID. This work comprehensively refactored the previous method while retaining the core module CSCV and made substantial improvements in multiple aspects, such as sports field initialization, iterative optimizer, and training methods, reducing the core indicator MID error from 48.9 to 38.44. The improved version significantly improved performance on various test benchmarks. The new contributions can be summarized as follows: (1) We introduce a self-supervised and truth mixed training pipeline that helps the network better balance the learning between occlusion fitting and texture matching, alleviate potential overfitting problems during training, and better predict motion prospects. (2) In order to solve the problem of standard GRU iteration modules having difficulty perceiving global motion, we propose an iterative optimizer GIR with a global perspective, which significantly improves the background performance of the network and slightly enhances the foreground performance while keeping the total time unchanged. (3) To alleviate the task alienation problem of the optimizer at different iteration stages, we constructed an initialization module to initialize the 3D flow field quickly. Reducing the total number of iterations while improving the optimizer’s optimization capability. (4) We further demonstrated the superiority of our method in the field of optical flow on the Sintel dataset. (5) Based on the self-supervised generalization method ADF, we provide a real-world generalization model for readers to test. Our code and test model can be obtained on https://github.com/HanLingsgjk/CSCV.
@@ II-A Optical Flow
Early methods[36, 37] viewed optical flow as an optimization problem for minimizing energy. Based on the regularization term and the constant brightness assumption, the optical flow results are optimized iteratively. Some recent excellent works have inherited their ideas[22, 38, 39], gradually updating the optical flow field through correlation matching and warping techniques. Compared with the previous methods [40, 23, 41, 42, 43, 44] of directly using convolutional networks to regress optical flow, the method based on correlation matching has a clearer theoretical background and more robust results. In this work, inspired by the successful experience of correlation matching, we extended this matching pattern to the scale dimension based on CSCV, improving the scale robustness of the optical flow baseline model.
It is worth mentioning that although some previous optical flow works[45, 46, 22] also used the idea of scale, our method differs fundamentally from theirs. Previous works are often based on multi-scale technology to increase the network receptive field, improve the inference speed, and strengthen the loss function. However, it is still image-to-image matching in essence. CSCV introduces a pattern for optical flow matching in scale space, a brand-new attempt.
@@ II-B Scene Flow
Methods Based on Point Cloud. Point cloud based methods[50, 51, 52, 53] capture 3D point clouds of the surrounding environment through an active laser or stereo camera, and establish the corresponding relationship between two frame point clouds to estimate 3D motion. However, this kind of method is limited by the sparse results and limited range, which is challenging to apply in more common video scenes. There is also a point cloud method based on the fusion framework, Camliflow[12]. By highly fusing the point cloud and the optical flow pipeline, the two pipelines can obtain complementary results at the same time. Camliflow achieves end-to-end scene flow estimation, but it still has the problem of depth information dependence.
Figure 3: Overview of ScaleFlow++. The network is mainly divided into two stages: initialization and iterative optimization. In the initialization stage, we construct a 4D correlation volume based on the 1/16 features extracted from ResNet and sample it using an all-zero optical flow field. Based on this, we regress the initialized 3D motion field. In the iterative optimization phase, we sample cross-scale correlation features from CSCV based on the current motion field, encode them, and send them to the GIR module for optimization. Output the final optimization result after N iterations.
II-C Motion-in-depth (MID)
MID describes the motion in the depth direction through the scale change of depth. It has many practical application scenarios, such as time-to-collision (TTC) estimation, LiDAR scene flow, scene flow, Etc. MID is also the core indicator of our ScaleFlow++.
Early TTC works[16, 17, 18] estimated the MID by tracking the motion trajectory of the interest points and building a motion model. Because the distribution of the interest points is random, this kind of method often makes it hard to pay attention to the primary target, yielding only sparse and low-precision results. Binary TTC[21] is a successful TTC scheme proposed recently, which can regress a dense MID field segmentally. Binary TTC sends the scaled image to the encoder to directly return the MID result of a certain range, and the specific range depends on the scaling factor. Therefore, a complete MID estimation requires multiple regressions to cover the complete MID interval. Unlike them, ScaleFlow++ only needs one calculation to get complete and dense MID results.
Yang & Ramanan[20] proposed an other monocular normalized scene flow method. They obtained dense MID results based on optical flow step-wise regression optical expansion and MID. However, the same as previous TTC work, they separated optical flow and MID estimation, resulting in the depth motion estimation result dependent on the optical flow result. Unlike them, ScaleFlow++ integrates optical flow and depth motion estimation into a harmonious task based on CSCV, making information blend and help each other. In addition, the cross-scale matching mechanism can also help the optical flow module more accurately estimate the objects with scale changes.
II-D Cross-scale Matching
Cross-scale matching aims to establish the correspondence between objects in different frames, and the scales of objects in these frames often have large changes, which is a difficult problem faced by current optical flow methods. In the early works, SIFT[19] obtains scale-invariant features in the scale space through hand-designed feature extraction operators, and then matches objects of different scales based on these stable features. In addition to being used for matching, cross-scale matching technology can also be used to estimate the scale changes of objects between frames. As early as 2005, Mikolajczyk[29] used Lindeberg’s scale selection method[28] to successfully estimate the optical expansion of clock towers captured at different focal lengths. Inspired by the above works, CSCV performs optical flow matching in scale space and estimates a dense scale change field based on the cross-scale matching results.
@@ II-D Cross-scale Matching
III Method
The framework of ScaleFlow++ is shown in Fig. 3, similar to RAFT, which constructs a Cross-Scale Correlation Volume (CSCV) on all pixel pairs and iteratively optimizes the 3D motion field through an GIR optimizer. We also constructed a small 4D correlation volume based on features of 1/16 size for rapid initialization of optical flow and scale flow, thereby reducing the optimization difficulty of the GIR. In the inference process, GIR directly outputs the optical flow $(f_{1},f_{2})$ and scale change $f_{3}$ fields for CSCV sampling. Combining previous work[20], we convert $f_{3}$ into motion-in-depth, which is convenient for training and can be more widely applied to downstream tasks.
Below, we will provide a detailed introduction to the construction and training process of ScaleFlow++, as well as how to convert network estimation results into a 3D motion field.